OpenStudy (anonymous):

Algebra Help Please! 1. Use the Zero Product Property to solve the equation. (x-9) (x-8) = 0 A) x = -8 or -9 B) x = 8 or 9 C) x = -2 or -3 D) x = 0 2. Use the Zero Product to solve the equation. (4k + 5) (k + 7) = 0 A) k = 5/4 or 7 B) k = 5/4 or -7 C) k = -5/4 or -7 D) k = -5/4 or 7 3. Solve by factoring. g^2 + 4g - 32 = 0 A) g = 32 or 28 B) g = 4 or 8 C) g = 2 or -16 D) g = 4 or -8 4. Solve by factoring. s^2 - 14s + 45 = 0 A) s = 3 or 15 B) s = -5 or -9 C) s = 5 or -9 D) s = 5 or 9 5. Solve by factoring. 2z^2 - 21z - 36 = 0 A) z = 6 or -6

OpenStudy (anonymous):

5. Solve by factoring. 2z^2 - 21z - 36 = 0 A) z = 6 or -6 B) z = 3 or -24 C) z= -3 or 24 D) z = -3/2 or 12

OpenStudy (anonymous):

@phi

OpenStudy (phi):

how far did you get ?

OpenStudy (phi):

btw, you should make only one question per post, otherwise it is too confusing

OpenStudy (phi):

if A is zero, do you know what is A*B ?

OpenStudy (anonymous):

@phi I don't understand what to do for any of the questions

OpenStudy (anonymous):

@Directrix

OpenStudy (phi):

The point is to learn the idea, and we start by knowing what is 0*1 ?

Directrix (directrix):

The Zero Product Property simply states that if ab = 0, then either a = 0 or b = 0 (or both). A product of factors is zero if and only if one or more of the factors is zero.

Directrix (directrix):

s^2 - 14s + 45 = 0 What are two numbers that multiply to 45 and add to -14 ? @Answers101

Directrix (directrix):

We need those numbers to factor unless you already know how s^2 - 14s + 45 factors.